### v2

# ]corresponding author

Abstract
The synthesis of metallic nanoparticle assemblies is nowadays well-controlled, such that these systems offer the possibility of controlling light at a sub-wavelength scale, thanks, in parts to surface plasmons. Determining the energy dispersion of plasmons likely to couple to light in these nanostructures is, therefore, a necessary preliminary task on the way to understanding both their photonic properties and their physical nature, in particular the role of the quadrupole contribution. Starting with a general model that takes account of all energy modes, we show that its low-lying energy dispersion, gained numerically, can be compared to that of a minimal model that treats dipoles and quadrupoles on the same footing. The main advantage of the latter relies on the fact that its formulation is tractable, such that a semi-analytical Bogoliubov transformation allows one to access the experimentally relevant energy bands. Based on this semi-analytical derivation, we determine quantitatively the limit of validity of both the dipole-only model and the presently proposed dipole and quadrupole model, compared to a full-plasmon-mode Hamiltonian. The results show that the dispersion relation, which accounts for dipoles and quadrupoles, is sufficient to capture the low-energy physics in most experimental situations. Besides, we show that at small lattice spacing, the contribution of quadrupoles is dominant around the Brillouin zone center.
I Introduction
Plasmonic materials have proven to show a wide range of properties, with continuous research ranging from the fields of photonics [1], condensed matter [2], and spectroscopy. As a few striking examples, quantum entanglement was shown to survive a photon-plasmon-photon conversion [3, 4], which justified a quantum description of the plasmons. A regime of very strong coupling with light was observed in plasmonic crystals, with the breakdown of the Purcell effect [5]. The possibility of investigating single-molecule vibrational states via plasmonic cavities was demonstrated [6, 7]. Among plasmonic materials, ordered plasmonic nanostructures are especially interesting as the control of their geometry enables the control of their optical properties [8, 9], and could lead to tailored properties of materials [10]. In the theoretical description of these superlattices, different models have been proposed. Some models use a classical description of the plasmons via point dipole [11, 12, 13] or point multipole models [14], and some use a quantum description of the plasmons via dipolar second quantization bosonic operators [15, 16, 17, 18] or multi-polar bosonic operators [19, 20, 21, 22]. The second quantized formalism becomes quite handy in describing the coupling with other fundamentally quantum systems, like molecules [23] or quantum dots [24] which makes it the formalism of choice here.
Regardless of the model chosen, it is always desirable to have a minimal model for the description of the physics at hand. That is why limiting the description to a coupled dipole model is often the chosen option. However, it is well established that this model stands for particles distant enough from one another. Indeed, a dipole-only model fails to describe the physics when the center-to-center distances ($d$) gets smaller than three times the radius ($d/R<3$) of the nanoparticles [25]. This is due to plasmon-plasmon induced hybridization [26], and it is observed even for small particles described in the quasistatic limit. This fact is particularly acute in recent interesting experimental realizations of colloidal crystals and other nanoparticle assemblies [5, 7], for which the nanoparticle center-to-center distance is well below the dipole model limit. Besides, considering coupling to fast electrons and their near fields requires higher order plasmons even for small particles, down to 4nm [27]. All these facts combine to suggest that a simple analytical expression of the low-energy plasmonic bands is necessary to help interpret future experimental data.
In the present paper, we intend to take stock of this knowledge by building a minimal semi-analytic description, in line with earlier investigations. This description will necessarily go beyond a simple dipole model, as we have previously emphasized. We thus investigate the importance of higher order plasmonic modes in dense assemblies. For the sake of clarity and simplicity, we choose a 1D chain of spherical nanoparticles as a model system, and we propose a minimal Hamiltonian leading to analytic expressions for the lowest energy plasmonic modes. Specifically, we discuss our derivation in light of previous similar approaches and show that taking account of the quadrupoles is sufficient, as it accurately describes the physics relevant to experiments. The present minimal model implies tractable Hamiltonian matrices, and therefore combines two advantages: it allows one to derive a semi-analytical expression of the low-lying excitation dispersion, and the contribution of dipoles vs quadrupoles can be easily investigated. Besides, it sets the stage for a future coupling to photons, which is more straightforward in contrast to full multipole models.
The paper is organized as follows: In the first section, we present the model as a truncation of the multipole model described in Refs.[25, 28], and show that it is comparable to Barros et al. approximation [21]. By inspecting the physical nature of the terms in the expression of the dispersion, we briefly argue why it is not possible to simplify this model further. Then, we present the plasmonic band structure obtained through this method and comment on the limit of validity. The paper closes with our conclusion and perspectives.
II Model
The optical properties of nanoparticle assemblies are formally described by the following generic Hamiltonian
$\mathcal{H}=H_{pl}+H_{\gamma}+H_{pl-\gamma}$ (1)
where $H_{pl}$ stands for the plasmonic degrees of freedom, whereas $H_{\gamma}$ and $H_{pl-\gamma}$ respectively describe the photonic bath and its coupling to the plasmons. For the sake of clarity, this paper focuses solely on determining the minimal $H_{pl}$ relevant to experiments. Indeed, the derivation of reliable physical observables first implies correctly describing the plasmonic Hamiltonian.
Figure 1: Schematic representation of a 1D chain of spherical particles of radius $R$ with center-to-center distance $d$. The chain runs along the $z$ axis.
In the present work, we consider an assembly of metallic spherical nanoparticles embedded in a dielectric matrix as depicted, in Fig.1, for a 1D case. In this system, each nanoparticle hosts a discrete spectrum of localized surface plasmons $\left|\phi_{\ell,m}\right\rangle$ with corresponding energies $\hbar\omega_{\ell}$, where $\ell$ stands for the order of the mode and $m$ stands for its azimuthal index. In second quantization, one can thus write a Hamiltonian $\sum_{\ell,m}\hbar\omega_{\ell}b_{\ell m}^{\dagger}b_{\ell m}$, where $b_{\ell m}^{\dagger}$ (respectively $b_{\ell m}$) creates (annihilates) a localized surface plasmon. For spherical particles, in the quasistatic limit, one has $\omega_{\ell}=\sqrt{\frac{\ell}{2\ell+1}}\omega_{p}$, where $\omega_{p}$ stands for the plasma frequency of the metal. To give an order of magnitude, from experimental values, the bulk plasmon frequency of silver is $\hbar\omega_{p}^{{\rm Ag}}=9.04{\rm\ eV}$ and the one of gold is $\hbar\omega_{p}^{{\rm Au}}=8.89{\rm\ eV}$ [29]. These relatively high-energy values translate to the important role played in plasmonics by the nanoparticles’ size, as well as the surrounding matrix in assemblies [30].
Within an assembly, the plasmon modes interact with one another, implying a hybridization, thereby forming bands. The plasmon-plasmon interaction terms can be evaluated either from overlap integrals $Q_{\ell m\ell^{\prime}m^{\prime}}$ [28, 31, 25] or from the Coulomb interaction in hydrodynamics models [26]. Taking this interaction into account and using the overlap formulation the plasmonic Hamiltonian reads
$\begin{array}[]{ccc}H_{pl}&=&{\displaystyle\sum_{\mathbf{k}}\sum_{\ell,m}}\hbar\omega_{\ell}b_{\mathbf{k}\ell m}^{\dagger}b_{\mathbf{k}\ell m}+{\displaystyle\sum_{\mathbf{k}}\frac{\hbar\omega_{p}^{2}}{4}\sum_{\ell,m}\sum_{\ell^{\prime},m^{\prime}}\frac{Q_{\mathbf{k}\ell m\ell^{\prime}m^{\prime}}}{\sqrt{\omega_{\ell}\omega_{\ell^{\prime}}}}}\left[b_{\mathbf{-k}\ell m}\left(b_{\mathbf{k}\ell^{\prime}m^{\prime}}+b_{\mathbf{-k}\ell^{\prime}m^{\prime}}^{\dagger}\right)+b_{\mathbf{k}\ell m}^{\dagger}\left(b_{\mathbf{k}\ell^{\prime}m^{\prime}}+b_{\mathbf{-k}\ell^{\prime}m^{\prime}}^{\dagger}\right)\right]\end{array},$ (2)
with the overlap for a 1D-chain given below, its general expression for other geometries can be found in Ref. [25]
$\begin{array}[]{lll}Q_{\mathbf{k}\ell m\ell^{\prime}m^{\prime}}&=&\left(\frac{R}{d}\right)^{\ell+\ell^{\prime}+1}\left(-1\right)^{\ell^{\prime}+m}\sqrt{\frac{\ell\ell^{\prime}}{(2\ell+1)(2\ell^{\prime}+1)}}\frac{(\ell+\ell^{\prime})!}{\sqrt{(\ell+m)!(\ell^{\prime}+m)!(\ell-m)!(\ell^{\prime}-m)!}}\\ &&\times\left[\text{Li}_{\ell+\ell^{\prime}+1}\left(e^{ikd}\right)+\left(-1\right){}^{\ell+\ell^{\prime}}\text{Li}_{\ell+\ell^{\prime}+1}\left(e^{-ikd}\right)\right]\delta_{mm^{\prime}}\end{array}$ (5)
and the Polylog function
$\text{Li}_{n}(z)=\sum_{k=0}^{\infty}\frac{z^{k}}{k^{n}}$ (6)
The main advantage of this formulation relies on the fact that, in its general expression, the Eq. (2) allows one to describe any plasmonic nanoparticle assembly within the quasistatic limit, provided that the corresponding overlap term $Q_{\bm{k},\ell m\ell^{\prime}m^{\prime}}$ is correctly evaluated. To avoid unnecessary cumbersome expressions that would obscure the discussion, the present work limits its quantitative investigation to 1D systems. Nevertheless, 2D and 3D situations can be addressed within the same framework. Yet, care has to be taken while computing the overlap of Eq. (5) from the real space expression in [25].
Owing to the symmetry of the chain, the present hamiltonian is block diagonal in $m$, and modes with different azimuthal indexes do not interact. In principle, as shown by Park and Stroud [25] a sharp description of the whole physics at play requires a large number of modes. However, if one focuses on the low-energy part of the spectrum, truncating at $\ell_{max}=2$ is sufficient, as shall be checked in Section III.1. In addition, from an experimental point of view this approximation appears reasonable: For instance, the plasmonic modes that will easily couple to light are $\ell=1$-dipolar modes and, as we shall see, those are strongly influenced by $\ell=2$-quadrupolar modes. Limiting oneself to $\ell_{max}=1$ in Eq. (2) leads to
$H_{dip}=$ $\sum_{\mathbf{k}}\sum_{m}\left\{\hbar\omega_{1}b_{\mathbf{k}1m}^{\dagger}b_{\mathbf{k}1m}\right$.
$\left.+\frac{\hbar\omega_{p}^{2}}{2\omega_{1}}Q_{\mathbf{k}11m}\left[b_{\mathbf{-k}1m}\left(b_{\mathbf{k}1m}+b_{\mathbf{-k}1m}^{\dagger}\right)+{\rm h.c.}\right]\right\}$ (7)
This latter Hamiltonian restricted to dipolar modes is investigated by Allard and Weick in Ref.[15]. In the same spirit, but keeping in addition quadrupolar modes, we shall truncate the Hamiltonian in Eq. (2) at order $\ell_{max}=2$. This approximation corresponds to that of Ref.[21].
Yet, the derivation of the Hamiltonian by Barros et al. is slightly different from the present one. It provides expressions of canonically conjugate momenta for quadrupolar excitations allowing for the coupling with a photonic bath, in the form of the quadrupolar tensor, and the expression is more involved. Nevertheless, in the quasistatic limit, the two models come together and provide access to the same physics. Besides, the present formulation allows one to investigate the role of higher-order multipoles, when Ref. [21], is limited to quadrupoles due once again to the difficulty of working out the coupling to free space electromagnetic field.
The diagonalization of dipole+quadrupole Hamiltonian resulting from the $\ell_{max}=2$-truncation of Eq. (2) is then gained by performing a Bogoliubov transformation. This unitary transformation is obtained once one has introduced the operators $\mu_{\mathbf{k}\ell}^{\dagger}$ and $\mu_{\mathbf{k}\ell}$ as linear combination of the original annihilation and creation operators $b$ and $b^{\dagger}$. Such that
$\mu_{\mathbf{k}\ell}=u_{\mathbf{k}\ell}b_{\mathbf{k}\ell}+v_{\mathbf{k\ell}}b_{\mathbf{-k}\ell}^{\dagger}+\sum_{\ell^{\prime}\neq\ell}\left(w_{\mathbf{k}\ell^{\prime}}b_{\mathbf{k}\ell^{\prime}}+x_{\mathbf{k}\ell^{\prime}}b_{\mathbf{-k}\ell^{\prime}}^{\dagger}\right)$ (8)
where $u_{\mathbf{{k\ell}}},$ $v_{\mathbf{{k\ell}}},$ $w_{\mathbf{{k\ell^{\prime}}}}$ and $x_{\mathbf{{k\ell^{\prime}}}}$ represent a set of complex coefficients which are nothing but the elements of the unitary matrix involved in the Bogoliubov transformation. Their expression do not enter the computation of the energies conducted below from the commutators expressions. This expression with $\left(\ell,\ell^{\prime}\right)=\left(1,2\right)$, enables to rewrite the Hamiltonian as
$H_{pl}=\sum_{\mathbf{k}\in\mathcal{D}}\sum_{\ell}\hbar\omega_{\mathbf{k}\ell}\mu_{\mathbf{k}\ell}^{\dagger}\mu_{\mathbf{k}\ell}.$ (9)
The sum in the reciprocal space runs over the half-space $\mathcal{D}$. Indeed, one has to pay attention to the fact that in the definition of Eq. (8) the original bosonic operators carry either a positive or a negative $k$. The expressions for the commutator $\mu_{\mathbf{k}\ell},H_{pl}]$ lead to solving $\det(M-\omega_{k}\mathds{1})=0$ with
$M=\begin{pmatrix}\omega_{1}+f_{\mathbf{k}11m}&-f_{\mathbf{k}11m}&f_{\mathbf{k}12m}^{*}&f_{\mathbf{k}12m}\\ -f_{\mathbf{k}11m}&\omega_{1}+f_{\mathbf{k}11m}&f_{\mathbf{k}12m}&f_{\mathbf{k}12m}^{*}\\ f_{\mathbf{k}12m}&f_{\mathbf{k}12m}^{*}&\omega_{2}+f_{\mathbf{k}22m}&-f_{\mathbf{k}22m}\\ f_{\mathbf{k}12m}^{*}&f_{\mathbf{k}12m}&-f_{\mathbf{k}22m}&\omega_{2}+f_{\mathbf{k}22m}\end{pmatrix}$ (10)
where $f_{\mathbf{k}\ell\ell^{\prime}m}$ is introduced
$f_{\mathbf{k}\ell\ell^{\prime}m}=\frac{\omega_{p}^{2}}{4}\frac{Q_{\mathbf{k}\ell\ell^{\prime}m}+Q_{\mathbf{-k}\ell^{\prime}\ell m}}{\sqrt{\omega_{\ell}\omega_{\ell^{\prime}}}}$ (11)
The low-lying energy dispersion across the Brillouin zone $\hbar\omega_{k}$ is given by
$\omega_{k}^{2}$ $=\frac{1}{2}\left(\omega_{1}^{2}+\omega_{p}^{2}Q_{\mathbf{k}11m}+\omega_{2}^{2}+\omega_{p}^{2}Q_{\mathbf{k}22m}\right)$
$-\frac{1}{2}\sqrt{\left(\omega_{1}^{2}+\omega_{p}^{2}Q_{\mathbf{k}11m}-\omega_{2}^{2}-\omega_{p}^{2}Q_{\mathbf{k}22m}\right)^{2}+4\omega_{p}^{4}|Q_{\mathbf{k}12m}|^{2}}$ (12)
The inspection of Eqs. (5) and (12) shows that the energy contains: dipole-dipole interaction terms $Q_{11}\propto\left(R/d\right)^{3}$, dipole-quadrupole interaction terms $Q_{12}\propto\left(R/d\right)^{4}$, and quadrupole-quadrupole interaction terms $Q_{22}\propto\left(R/d\right)^{5}$, each of which having a non-trivial dependence in $k$.
Two interacting subspaces enter Eq. (12): that of dipoles and that of quadrupoles. This observation leads to the rewriting of the dispersion as $\omega_{k}^{2}=\frac{1}{2}(\omega_{1k}^{2}+\omega_{2k}^{2})-\frac{1}{2}\sqrt{\left(\omega_{1k}^{2}-\omega_{2k}^{2}\right)^{2}+4f_{12k}^{2}}$. Besides, the analysis of the amplitudes of the respective elements $\omega_{1k}$, $\omega_{2k}$ and $f_{12k}$ over the Brillouin zone shows that $f_{12k}^{2}\ll\omega_{1k}^{2},\omega_{2k}^{2}$ as can be seen in Fig. 2. Therefore, it is tempting to perform a perturbative expansion of Eq. (12) with respect to the parameter $\eta$
$\eta(k)=\frac{|Q_{\mathbf{k}12m}|}{|\omega_{1}^{2}+\omega_{p}^{2}Q_{\mathbf{k}11m}-\omega_{2}^{2}-\omega_{p}^{2}Q_{\mathbf{k}22m}|}$ (13)
Figure 2: Terms appearing in the dispersion relation of the model with schematic descriptions of the surface charge densities associated to the $k=0$ and $k=\pi/d$ modes. The dotted blue line corresponds to interacting dipoles, while the orange solid line corresponds to a chain of interacting quadrupoles. As pointed out in the text, the bands cross for dense enough chains, here the values correspond to $d/R=2.4$ and $m=1$.
This would correspond to a downfolding of the two subspaces previously mentioned to an effective one corresponding to dipoles dressed by quadrupolar excitations. However, one has to notice in Fig. 2 that for dense chains with $d/R<3$, the dipolar band $\omega_{1k}$ and the quadrupolar band $\omega_{2k}$ cross as one goes from $k=0$ to $\pi/d$. Consequently $\eta$ diverges at this crossing and the nature of the low-energy mode changes. All these facts together suggest that the minimal model required to describe the low-energy excitations in an assembly of plasmonic nanoparticles is given by the Hamiltonian of Eq. (2) with $\ell_{max}=2$. To support these words, we use the dispersion relation (12) for different chain densities and different plasmon polarizations (i.e., different $m$) and compare it to other models. In the following section we consider chain densities relevant to experiments and outside of the dipole model validity domain.
III Results and discussion
III.1 Superlattice spacing $d$
Experimentally, dense metallic nanoparticle assemblies can be synthesized as colloidal crystals formed with small ligands [5], with inter-particle distances down to $1\sim 4\ {\rm nm}$. There has been remarkable success in controlling the lattice parameter for tunable crystals with DNA strands [32] or polymer shells [33]. These experimental realizations generally give rise to situations for which $2.4\lesssim d/R\lesssim 3$. Denser assemblies with $1\ {\rm nm}$ particle separations or less are also of experimental relevance. However, in that case, tunneling effects occur [34], and one has to resort to a full quantum description, which is beyond the scope of the present work. Having that in mind, the quantitative analysis described below involves three distances: $d/R=3$, $d/R=2.6$ and $d/R=2.4$.
The plasmon-plasmon interaction part of the Hamiltonian in Eq. (2) contains a prefactor, explicitly written in Eq. (5), with an algebraic decay $d^{-(\ell+\ell^{\prime}+1)}$. This describes how the superlattice spacing influences the relative importance of the plasmonic multipoles, with higher-order modes gaining more importance as the particles get closer together. Nevertheless, the experimentally relevant modes are those likely to couple to light, namely and principally dipolar ones. Therefore, it becomes essential to understand qualitatively and quantitatively the role of these higher-order modes on the lower energy plasmonic band, which contains dipolar modes.
The dispersion of the different modes across the Brillouin zone is shown in Fig. 3 for each case: $d/R=3$, $d/R=2.6$, and $d/R=2.4$. For the sake of simplicity, one can take a formal example of $R=10{\rm\ nm}$, which is relevant to the quasistatic limit and provides interparticle distances of 24 nm in the $d/R=2.4$ case, in line with some typical colloidal crystals [32]. In fact, as the model is scale invariant, its limitation depends mainly on the quasistatic approximation validity. The converged $\ell_{max}=20$-model is plotted as gray and solid black lines, whilst the red dashed lines correspond to the dipole-model, and the blue crosses to the dipole+quadrupole model. The two low-lying highlighted bands are directly linked to dipolar modes. This is obvious if one looks at the $d/R=3$ case, where the three models produce nearly the same dispersion, and the deviation from a pure dipole model gets more pregnant as the inter-particle distance gets smaller. To understand the underlying mechanism, it is necessary to determine the nature of each band.
Figure 3: Plasmonic band diagram of a chain of spherical metallic nanoparticles of radius $R=10$ nm with three different lattice spacing (a) $d=30\ {\rm nm}$, (b) $d=26\ {\rm nm}$ and (c) $d=24\ {\rm nm}$. Black lines are calculated with $\ell_{max}=20$, red dashed lines are the dipole model eigenvalues with $\ell_{max}=1$, and the blue crosses are the $\ell_{max}=2$ model taking into account quadrupoles. The orange solid lines in (c) correspond to $\ell_{max}=3$ for comparison. In (a), (b) and (c) the high energy bands calculated with $\ell_{max}=20$ are grayed out for readability. Insets show the corresponding chain densities.
The two highlighted bands in Fig. 3 correspond to the $m=0$ and $m=\pm 1$ indexes, the latter being two times degenerate owing to the symmetry of the chain. In the dipole model, the $m=0$ mode corresponds to dipoles aligned along the chain, while the $m=\pm 1$ relates to transverse modes with dipoles orthogonal to the chain axis. At the Brillouin zone center, the $m=0$-mode with all dipoles pointing in the same direction is the fundamental excitation, as it minimizes the dipolar tensor and can be viewed as a bonding mode. In contrast, at the edge $k=\pi/d$, a Néel-like order for the $m=1$-mode is lower in energy. This $\pi/d$ mode is referred to as an anti-bonding mode in the hybridization picture proposed by Nordlander et al. [26]. Therefore, a band-crossing occurs as one goes from $k=0$ to $\pi/d$, and this general observation does not depend on the superlattice spacing. However, the dipole model remains quantitatively valid only for $d/R\gtrsim 3$ but at smaller inter-particle spacings and finite $k$ this simple approach fails to provide an accurate description. The blue dashed line in Fig. 4 highlights this fact: it represents the relative difference $\Delta(k)=\left|\left(\omega(k)-\omega_{\ell=20}(k)\right)/\omega_{\ell=20}(k)\right|$ with respect to the converged $\ell_{max}=20$-model. In addition, this is further confirmed by the computation of the coefficient of determination $\mathsf{R}^{2}$
$\mathsf{R}^{2}=1-\frac{\sum_{k}(\omega(k)-\omega_{\ell=20}(k))^{2}}{\sum_{k}(\omega(k)-\overline{\omega_{\ell=20}})^{2}},$ (14)
which is given in Table 1.
$\ell_{max}=1$ $\ell_{max}=2$
$m=0$ $m=\pm 1$ $m=0$ $m=\pm 1$
$d/R=3.0$ 0.98 0.95 1.00 1.00
$d/R=2.6$ 0.88 0.71 1.00 0.99
$d/R=2.4$ 0.57 0.36 0.96 0.95
Table 1: Values of $\mathsf{R}^{2}$ coefficient between the $\ell=20$ values and the truncated models values for the three lowest energy bands of different superlattices geometry corresponding to the plasmonic band diagrams of Fig. 2.
In contrast, the dipole+quadrupole model, depicted by the blue crosses in Fig. 3 remains quantitatively valid down to $d/R=2.4$. As one can see in Fig. 4, truncating at $\ell_{max}=2$ leads to a relative energy difference that remains below $3\%$ across the whole $k$-space, even for dense assemblies and an $\mathsf{R}^{2}$, which is larger than $95\%$, comparable to that of the dipole model at $d/R=3$. Taking into account the higher-order $\ell=3$ octupolar modes corrects the model prediction only slightly as can be seen in Fig. 3, with $\mathsf{R}^{2}$ values of 0.99 at $d/R=2.4$ for both bands. In principle, the exact solution implies a Hamiltonian diagonalization with $\ell_{max}\to\infty$, it is therefore a matter of taste to truncate at an arbitrary mode once a convergence criterion is satisfactory. Of course, involving the octupoles implies a better quantitative match, but limiting ourselves to a dipole+quadrupole truncation has the advantage of providing a good physical description of the main low-lying excitations (within 5% quantitatively speaking). Besides, this approach allows one to obtain the tractable semi-analytical dispersion of Eq. (12), and provides a clear framework to understand the role of higher-order multipoles in plasmonic bands, as we shall see in the next Section.
As for the calculation of the plasmonic eigenvalues at a given $\bm{k}$, the dimensions of the Hamiltonian to be solved increase as $\sum_{\ell=1}^{\ell_{max}}(2\ell+1)=\ell_{max}^{2}+2\ell_{max}$. This gives Hamiltonian sizes of 3, 8, and 440 for the dipolar model, the quadrupolar model, and the model with $\ell_{max}=20$ respectively. From an experimental point of view, the main interest relies in the high symmetry $k=0$ point. To quantify the energy difference between the dipole model prediction, and the dipole+quadrupole model at this point, let us consider silver nanoparticles with $\hbar\omega_{p}=9.04{\rm\ eV}$. The dipole model’s predictions differ from those of the converged model $\ell_{max}=20$ by $\Delta E_{m=0}(k=0)=0.09{\rm\ eV}$ and $\Delta E_{m=1}(k=0)=0.29{\rm\ eV}$ while for the dipole+quadrupole model we have $\Delta E_{m=0}(k=0)=0.09{\rm\ eV}$ and $\Delta E_{m=1}(k=0)=0.05{\rm\ eV}$.
Figure 4: Relative energy differences $\Delta$ with regard to the converged model $\ell_{max}=20$ across the Brillouin zone for (a) $d/R=3$, (b) $d/R=2.6$ and (c) $d/R=2.4$ in the same order as in Fig.3 and for both $m=0$ and $m=\pm 1$. The expression for $\Delta$ is $\Delta=\left|\left(\omega(k)-\omega_{\ell=20}(k)\right)/\omega_{\ell=20}(k)\right|$.
The high precision that is obtained from the simple truncation at $\ell_{max}=2$ comes from the very structure of the Hamiltonian of Eqs. (2) and (5). Indeed, as was previously emphasized, only terms coupling the same $m$ remain; this gives rise to a sparse block-diagonal matrix, which combines to the decay in $d^{-\left(\ell+\ell^{\prime}+1\right)}$ to make the higher-order modes contribute very little. To sum up, these results confirm that the dispersion of Eq. (12) is sufficient to describe low-energy physics in plasmonic nanoparticle chains.
III.2 Eigenstate analysis: role of the quadrupoles
To determine more precisely the role of the quadrupoles in the system, let us inspect the eigenstates of the low-lying plasmonic bands. These eigenstates can be written as linear combinations of the single-particle multipoles
$\left|\psi_{m}(k)\right\rangle=a_{1m}(k)\left|\phi_{1m}\right\rangle+a_{2m}(k)\left|\phi_{2m}\right\rangle,$ (15)
where the expressions of the coefficients $a_{1m}(k)$ and $a_{2m}(k)$ depend on the ordering of the subspaces of dipoles $\omega_{1k}^{2}$ and quadrupoles $\omega_{2k}^{2}$ and their difference $\Delta_{m}(k)=\omega_{1}^{2}+\omega_{p}^{2}Q_{k11m}-\omega_{2}-\omega_{p}^{2}Q_{k22m}$, appearing in the square root of Eq. (12). This leads to the following rewriting
$\left\{\begin{array}[]{lll}\left|\psi_{m,>}(k)\right\rangle&=&-\sin\frac{\theta_{m}(k)}{2}\left|\phi_{1m}\right\rangle+\cos\frac{\theta_{m}(k)}{2}\left|\phi_{2m}\right\rangle\\ \\ \left|\psi_{m,<}(k)\right\rangle&=&\cos\frac{\theta_{m}(k)}{2}\left|\phi_{1m}\right\rangle-\sin\frac{\theta_{m}(k)}{2}\left|\phi_{2m}\right\rangle\end{array}\right$. (16)
where $\left|\psi_{m,>}(k)\right\rangle$ stands for the case $\Delta_{m}(k)>0$, whilst $\left|\psi_{m,<}(k)\right\rangle$ holds for $\Delta_{m}(k)<0$, and the phase $\theta_{m}(k)$ is given by $\theta_{m}(k)=-\tan^{-1}\left|\frac{2Q_{k12m}}{\Delta_{m}(k)}\right|$.
Fig. 2 ($d/R=2.4$) shows that the sign of $\Delta_{m}(k)$ might change along the Brillouin-zone, implying a crossing of dipolar and quadrupolar subspaces. Therefore, as seen in Eqs. (16), the main component of the mode changes, switching from dipoles to quadrupoles or vice versa. If the lattice spacing is large enough ($d/R\gtrsim 3$), this crossing does not occur since quadrupoles are too high in energy and almost decoupled from dipoles. In contrast, when the particles are close enough, the admixture of the modes can no longer be neglected as their energies are comparable. This implies that even though one focuses on low-energy physics, quadrupoles must be considered in the model since they become a non-negligible component of the bands. To highlight this fact, we plot the $a_{1m}(k)$ and $a_{2m}(k)$ coefficients for the three different lattice spacings $d/R=3$, $d/R=2.6$ and $d/R=2.4$, and the two different bands $m=0$ and $m=\pm 1$, in Figs. 5 and 6.
For $d/R=3$, depicted at the top of Figs. 5 and 6, as expected, the dipolar component largely dominates the wave-function of Eq. (16) across the whole Brillouin zone. In the case of the $m=0$ band, and for both $d/R=2.6$ and $d/R=2.4$, the wave-function is purely dipolar at $k=0$, then it keeps a major dipolar contribution for small wave-vectors. At larger ones, a crossing occurs, and a pure quadrupolar state ends to take place at $k=\pi/d$. At high-symmetry points, quadrupoles and dipoles do not couple, which is in line with the results of Barros et al. [21], they only mix at finite wave-vectors and eventually cross. Even though the crossing wave-vector gets smaller as the lattice spacing is reduced, it remains in the order of $k_{{\rm cross}}\sim 0.6\pi/d$, which means that this would hardly couple to light.
More interestingly, for the $m=\pm 1$ bands of Fig. 6, the nature of the state at $k=0$ changes as the lattice spacing is reduced: From a pure dipolar state for $d/R=3$, it becomes purely quadrupolar for $d/R=2.6$ and $d/R=2.4$. As $k$ is increased, the dipolar and quadrupolar components mix, up to the crossing point $k_{{\rm cross}}\simeq\left(0.1-0.2\right)\pi/d$ from which the dipolar state becomes dominant. In contrast to the previous $m=0$ case, the crossing point is close to the Brillouin zone center and, therefore, accessible to light-scattering.
As a result, it should be experimentally possible to measure the relative contribution of each component, dipole and quadrupole, and to see the physical nature of the state change if the assembly is more or less dense. However, this work of coupling to light and a photonic bath implies a modification of the eigenstates and is left for further investigation.
Figure 5: (a) Components of the eigenfunctions associated with the band dispersion (12) for the $m=0$ band of a chain of spherical nanoparticles of radius R=10R=10nm with three different lattice spacing (a) $d=30nm$, (b) d=26d=26nm and (c) d=24d=24nm. Red dashed lines are the dipole component $a_{\ell=1,m=0}(k)$, and the blue lines are the quadrupolar components $a_{\ell=2,m=0}(k)$
Figure 6: (a) Components of the eigenfunctions associated with the band dispersion (12) for the $m=\pm 1$ band of a chain of spherical nanoparticles of radius R=10R=10nm with three different lattice spacing (a) $d=30nm$, (b) d=26d=26nm and (c) d=24d=24nm. Red dashed lines are the dipole component $a_{\ell=1,m=1}(k)$, and the blue lines are the quadrupolar components $a_{\ell=2,m=1}(k)$
IV Conclusion and perspectives
Having developed a quasistatic Hamiltonian (2) from the plasmonic overlap of classical and the quantum models for plasmonic arrays, we have shown that dipolar and quadrupolar modes represent the major contribution of the low-energy physics and can be decoupled from higher-order modes. Consequently, we proposed a minimal model that treats dipoles and quadrupoles on the same footing and allows an accurate description of the two first plasmonic excitations across the whole Brillouin zone of a chain of nanoparticles. Further inspection of the nature of the modes underlined the importance of quadrupoles in dense particle arrays, solely due to plasmon-plasmon interactions. For this reason, in future work considering also the coupling with light, the quadrupole will have to be accounted for regardless of how little they couple to light. We pointed out that even though quadrupoles do not couple to dipoles at high symmetry points ($k=0$ for the chain), they do at others, rendering them essential to investigate plasmon-polaritons in nanoparticle assemblies. The coupling of the present Hamiltonian to a photonic bath $H_{\gamma}$ that would allow taking account of retardation effects and computing physical observables is left for further investigation. Finally, the second quantized formalism used here makes our Hamiltonian relevant for coupling to other systems, such as molecules or quantum dots, that couple to the highly localized fields of quadrupoles.
Acknowledgments
This work is supported by the Labex SOLSTICE (ANR-10-LABX-0022)